OpenStudy (anonymous):

find the distance between the points (-5,7) and (4,7)

OpenStudy (anonymous):

the formula for computing the distance between 2 points is: d = sqrt[(x1-x2)^2 + (y1 - y2)^2] d = sqrt [(-4-5)^2 + (-7-6)^2] d = sqrt [(-9)^2 + (-13)^2] d = sqrt [ 81 + 169 ] d = sqrt (250) d = 15.811

OpenStudy (anonymous):

ah im confused

OpenStudy (anonymous):

She is explaining the distance formula to you :P

geerky42 (geerky42):

\[d = \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] Where \((x_1,y_1)\) and \((x_2,y_2)\) are given points.

geerky42 (geerky42):

In your case, you have (-5,7) and (4,7)

geerky42 (geerky42):

So after plugging in, you should have \(d = \sqrt{(-5-4)^2 + (7-7)^2}\)

geerky42 (geerky42):

Shortcut here is that since you can see that y-coordinates of points are same here. So this problem basically asks you to find the distance between -5 to 4.

geerky42 (geerky42):



OpenStudy (anonymous):

that 10 right

geerky42 (geerky42):

\(4-(-5) = 10\)?

OpenStudy (anonymous):

i suck at math so im not sure if im correct

OpenStudy (anonymous):

wait it 9

geerky42 (geerky42):

Correct. \(4-(-5) = 4+5 = \boxed 9\)

OpenStudy (anonymous):

yeah i did it

OpenStudy (jennyrlz):

two negatives make a positive :)

OpenStudy (anonymous):

thank u so much @geerky42

geerky42 (geerky42):

No problem